### v1

@@ 6.1.3 Learning strategies
In order to reduce gradient variance from different domains, Atwany and Yaqub [10] presented a method  [^18] https://github.com/BioMedIA-MBZUAI/DRGen that encourages seeking a flatter minima while imposing a regularization. Their method is inspired by Stochastic Weight Averaging Densely (SWAD) [30] and Fishr [152].
Inspired by the concept of Stochastic Weight Averaging (SWA) [86], SWAD seeks a flat minimum by averaging the weights "densely", i.e., by iterations (rather than by epochs). In addition, to avoid overfitting, the start and the end iterations for collecting the weights to be averaged are determined based on the validation loss. Specifically, the start iteration is defined as the iteration where the validation loss first reaches a local minimum while the end iteration is defined as the iteration when the validation loss no longer decreases but keeps increasing. This method enables averaging weights only from specific iterations where the validation loss decreases. It has shown promising performances for DG.
On the other hand, Fishr [152] is a regularization approach that enforces domain invariance in the space of the gradients of the loss. In particular, the domain-level variances of gradients are matched across training domains. This regularization enhances the out-of-distribution generalization properties by aligning the domain-level loss landscapes at convergence. It uses the gradient covariance, the Hessian of the loss and Fisher information relationship. Fishr constraints domain invariance to the classifier’s weights using an exponential moving average after a predetermined warm-up phase, where the network learns predictive features. Once the warm-up phase is finished, Fishr regularization enforces the domain-level gradient invariance in the classifier by aligning the gradient covariances at the domain level. The Fishr loss is thus formulated as follows:
@@
The total loss is then formulated as:
$\mathcal{L}_{Total}=\frac{1}{M}\sum_{i=1}^{M}R_{i}+\lambda\mathcal{L}_{Fishr}$ (21)
where $\lambda$ is a hyperparameter and $R_{i}$ denotes the average empirical risk averaged for all training domains.
Atwany and Yaqub [10] adopted Stochastic Weighted Domain Invariance where they used Fishr loss for constraining invariance based on the difference in covariance matrices. After the warm-up phase, that is when the Fishr regularization is activated, they used exponential moving average to compute stable gradient variances and applied a penalty anneal iteration coefficient to help moderate the update speed at each time-step. They also integrated in their framework stochastic weighted averaging.
@@
Bissoto et al. [23] proposed a DG approach based on DRO. This approach assume the presence of multi-domain data. To this end, their pipeline consisted of first preparing the data. They partitioned the data into biased training and test sets (named trap sets) to better assess model generalization. In this context, "trap set" refers to datasets whose training and test splits present high and opposite correlations between the annotated artifacts and target label (malignant vs. benign). Afterwards, the training set is divided into artifact-based domains. Data are grouped according to the presence of specific artifact, such as dark corners or drawing of rulers. In other words, each training source domain contains the same set of artifacts. Next, these training source domains are used to train a robust learning algorithm, the Group Distibutional Robust Optimization (GroupDRO) [160] algorithm, which consists of minimizing the loss of the worst-case training source environment. In the last phase, the authors employed a test-time debiasing procedure to reduce the influence of spurious features in the inference images. The experimental results for skin lesion detection showed that GroupDRO allows learning more robust features.
@@
To improve the model generalizability, Lin et al. [119]  [^19] https://github.com/linzhlalala/CVD-risk-based-on-retinal-fundus-images proposed a multi-task network for cardiovascular disease risk (CVD) estimation using fundus photographs. Inspired by the DG research and contrastive learning, they proposed a Cross-Laterality Feature Alignment (CLFA) for CVD estimation in fundus images. They hypothesized that fundus photographs share invariant representation for the CVD risk over the two eyes. A DL model identifying the invariant representation, would have a better generalization capability. Therefore, they proposed to adopt a Siamese network and pre-train it using the left and right fundus photographs for each patient as positive sample pairs. They added a supervised learning branch to their Siamese network where they jointly trained the model on WHO-CVD score and on seven clinical variables explicitly correlated with WHO-CVD such as age, systolic blood pressure, total cholesterol, body massive index, gender, smoking status and diabetes status. They also integrated a feature-level knowledge distillation. Given two input images (left and right fundus photographs) from a single patient, the feature with the smallest supervised learning loss is considered as the teacher whereas the other as the student. For the teacher-level features, they performed stop-gradient operation when updating the feature extractor. The results showed that CLFA pre-training reduced the feature-space discrepancy between the UK biobank dataset collected using the Topcon 3D OCT-1000 MKII and the other cameras (Mediwork portable camera).
@@
In contrast to the domain-invariant methods, Wang et al. [206] embraced the disharmony in data and developed a simple framework to cope with domain shift problem. The main idea is to design a pre-trained classifier using the source data and adapt it to new data. For adaptation, they studied scenarios where we do not have access to ground-truth label of the target domain. They used auxiliary task (predicting age, gender and race) and integrated them in their framework. Specifically, their method consisted of four stages: 1) pre-training a classifier (composed of a feature extractor, an auxiliary task network and a primary task network) using the source data on the primary task and the auxiliary labels, 2) adapting the feature extractor: fine-tuning the feature extractor and the auxiliary task network for the auxiliary tasks using data from the target domain while constraining the feature extractor not to change significantly, 3) adapting the primary task classifier, the primary classifier is fined-tuned on the source data to correctly classify the primary task (i.e., classification) based on the modified features, 4) inference, inference on data coming from the target domain is performed using the updated weights of the feature extractor and the primary classifier. The authors showed improvement of performances when tested on 3D brain MRI dataset for the classification task of Alzheimer’s disease and schizophrenia.
@@
For MIDOG challenge, Razavi et al. [155] proposed a multi-stage mitosis detection method based on a Cascade R-CNN. This method was developed to be sequentially more selective against false positives. The Cascade-RCNN comprises a sequential detectors with increasing intersection over union to reduce false positives which may be attributed to the hard to detect mitotic cells. It consists of two-stage: 1) a region proposal network that detects candidate region and 2) a region proposal network and a classification network that performs classification on the candidate regions. This method achieved a F1 score of 0.7492 on the MIDOG testing set. This ends our presentation of DG solutions to address covariate shift.
6.2 Concept shift and label noise in medical imaging
@@
The quality of annotations has a crucial role on the model generalizability, which is usually evaluated by applying the model to an independent test dataset. Nevertheless, the annotation process can be subjective, and the issue of label noise is sometimes unavoidable. In addition, the quality of annotations can differ among various annotators. To improve prediction performances, DL methods have been proposed to address the problem of noisy labels and concept shift. These methods can be categorized into two main classes: data-centric methods and model-centric methods. Data-centric methods focus on data adjustment and transformation (Section 6.2.1). These methods focus on identifying noisy samples and correcting their labels. Model-centric methods include learning methods (Section 6.2.2) and collaborative methods (Section 6.2.3). Learning methods propose an optimization framework based on loss functions (i.e., regularization), architecture (e.g., graphs) and learning strategies (active learning, zero-shot learning, meta-learning, etc.). On the other hand, collaborative methods exploit the cooperation between models to boost DL performances.
@@ 6.2.1 Data adjustment and transformation
Son et al. [182] proposed a method to filter mislabeled data based on discrepancies between the network’s logits and given labels. More precisely, their method uses the softmax outputs to judge the correctness of the given labels. After filtering suspected mislabeled data, they masked the loss function computed over such samples.
They evaluated their method for detecting referable diabetic retinopathy using Kaggle 2015 dataset. The dataset was split into three subsets: a training subset, a validation subset and a testing subset. The training set was corrupted by randomly flipping referable DR to non-referable and vice versa with probabilities ranging between 0 and 0.8. To detect mislabeled samples in the training stage, they introduced a filtration network that is trained on top of the classification network. More formally, the filtration network is trained to minimize the logistic regression loss over validation data. It is only trained at each validation step (one training epoch) in the training phase, it does not contribute in the testing stage. Since the filtration network is trained on clean labels, it is able to detect clean positive data by assigning high values on positive images with clean labels and assigning low values for negative images. The authors thus examine the classifier’s confidence to perform data filtration and reduce the effect of the incorrect annotations. The filtration network is used to mask the loss values for suspected noisy samples. They showed that their proposed strategy outperformed state-of-the art methods such as S-model and Bootstrap when noise ratio is 0.2.
@@
Xue et al. [218] used the principle of Confident Learning (CL) [139] to identify noisy samples in the training data. It estimates label uncertainty which is the joint distribution between the noisy and true labels. To this end, similar to the previously mentioned approach, it employs the predicted probability outputs from a classification DL model. In CL, the class imbalance and heterogeneity in predicted probability distributions across classes are addressed by using a per-class threshold when calculating the confident joint. To further enhance the precision, Xue et al. [218] proposed an ensemble strategy. It consists of training three different classification networks and selecting the candidates that were jointly identified as noisy using CL. The proposed approach was assessed for automated visual evaluation for precancer screening. It achieved good performances with a kappa score of 0.687 when using cleaned train and validation sets compared to a score if 0.682 using the original train and validation sets.
To address the problem of weak label, Uncertainty-Aware Sampling Framework (UASF) [^20] https://github.com/machiraju-lab/UA-CNN, was proposed by Aljuhani et al. [3]. Example of weak labels include diagnosis of WSI-level labels obtained from patient records, where there is minimal annotation indicating the specific regions within these images that justify the diagnosed label at the WSI-level. The UASF comprises two stages. The first stage comprises a screening stage, where a screening model, a Resnet18 backbone with modifications for uncertainty estimation via Monte Carlo dropout, is trained on all available training tiles to screen for informative tiles. Next, in order to identify the most relevant tiles, an informative sampling is employed. The second stage, an uncertainty-guided model, is trained on the refined training dataset.
The informative sampling algorithm operates on the principle that tiles most predictive of their WSI level label exhibit high prediction probability and low uncertainty. It utilizes a model’s output, combining prediction probabilities from the overall model and uncertainty measures derived from Monte Carlo inference, to identify these tiles. To achieve this, the algorithm first models uncertainty as a quadratic function of prediction confidence for each WSI. It aims to identify a selection threshold for identifying the most relevant tiles. To this end, a threshold is determined based on the overall uncertainty measure of the WSI samples and the variability of tile predictions. The algorithm computes the difference between the True Predicted Rate (TPR) and the False Predicted Rate (FPR) at each threshold level, selecting the minimum threshold value as optimal.
@@
In the area of genomics and computational biology, Bai et al. [14] [^21] https://github.com/BRF123/Cnngeno proposed a novel method for calling genotypes from next-generation sequencing data, specifically targeting deletions. Their method is called Cnngeno, it converts sequencing data into images for genotypes classification. It adopts a convolutional bootstrapping strategy to improve its ability to handle noisy labels in the data. First, the authors manually selected a set of highly reliable seeds (i.e., a subset of samples) as training set and trained their model until convergence. After the initial training, the model is used to classify the remaining samples in the dataset. The process involves selecting samples with higher classification confidence into a seed set, and this selection process is based on a set classification confidence level, which is set at 0.8. The expansion of the seed set is repeated until no new seeds are added, and the final trained model is then used for classifying genotypes.
Xu and Chen [213] tackled the problem of label noise in neural decoding. They proposed a deep neural network based on sample re-weighting method. In their work, they assume that the training set is noisy while the validation set is small but accurately labeled. The sample re-weighting strategy consists of assigning weights to training samples to minimize the loss on a clean unbiased validation dataset to solve the problem of label noise and improve prediction performance. The process involves adjusting the weights of the training samples in such a way that samples which are likely to be correctly labeled have a higher weight, while those that are likely mislabeled are down-weighted. The proposed method was assessed on simulated neural activity data and calcium imaging data of anterior lateral motor cortex. It has shown that it correctly predicted the behavioral variable even with 36% of samples in the training set are mislabeled in the simulated data. On the anterior lateral motor cortex study, the proposed method, which was trained with 48% of training mislabeled samples, can predict trial types with an F1 score of around 0.85.
Similar to Xu and Chen [213], to effectively alleviate the disturbance from data labeling errors, Hu et al. [80] adopted a sample re-weighting [^22] https://github.com/TwistedW/MIAV strategy, where higher weights are assigned to clean samples and lower weights to noisy samples. First, the output features of the encoder are grouped into $k$ groups (where $k$ was set to 2) according to the channel layer. Next, a linear projection is performed on the grouped results after global average pooling. Finally, they applied a group weighted learning on the results of the two groups after ReLu activation. They evaluated their approach on the automatic classification of retinal Arteries and Veins, which is essential for diagnosing cardiovascular and eye-related diseases. The proposed method showed good results achieving an accuracy of 97.47%, 96.91%, 97.79%, and 98.18% on AV-DRIVE, HRF, LES-AV and a private dataset, respectively. Therefore, the sample re-weighting strategy effectively alleviates the influence of noisy labels due to the subjective judgment of annotator.
@@
Hermoza et al. [78] [^23] https://github.com/renato145/CASurv proposed a method for predicting survival time from medical images using both censored and uncensored data. Censored data are cases where the event of interest (such as death) has not occurred by the end of the study. These data are typically ignored in previous studies, resulting in an under-utilization of the training data. To leverage these incomplete observations, the authors used the concept of pseudo labels, i.e., using the output of the network to estimate the label. For each censored data point, a pseudo label is estimated, which represents a lower-bound survival time. The authors argue that the quality of generated pseudo-labels depends on the training procedure stage: during the first epochs, the pseudo-labels are less accurate than the ones at the last epochs. Therefore, they proposed to control the ratio of censored sample labels to be replaced with pseudo labels at each epoch using a cosine annealing schedule. Using this strategy, no pseudo-labels are produced at the beginning of the training and by the end of the training all censored data with their pseudo labels are produced. The evaluation of their proposed method on pathology and X-ray images from the TCGAGM and NLST datasets showed good prediction survival accuracy on both datasets.
@@
Bai et al. [15] proposed a label correction method in a semi-supervised approach. Specifically, they generate new high quality labels from unlabeled data. They used the concept of adaptive threshold to reduce noisy labels and retain the effective information.Their method called Pseudo-Labeling based on Adaptive Threshold (PLAT) reduces the generation of noisy labels and therefore enhance the generalizability of the models. It is inspired by epistemic uncertainty [97]. Unlabeled images are inputted to the model $k$ times using Monte Carlo Dropout. This will result into $k$ predictions. Uncertainty of pixels is computed by calculating the variance of $k$ predictions. Finally, a normalization step is achieved by dividing the results with the largest variance among all predictions. The normalized result is used as an adaptive threshold. Compared to model trained only on labeled images, this method showed a gain of 9%-13% in terms of F1 score. These performances are comparable to a model trained using supervised learning paradigm.
@@
Qiu et al. [148] proposed a self-training strategy consisting of noisy label cleaning optimization. First, the model is pretrained using the initial noisy labels, with a large fixed learning rate. This strategy assumes that the network does not overfit the label noise and maintain high performance when using a high learning rate. Then, noise-free labels (soft labels) are initialized based on the pretrained model. These labels are computed using the softmax output of the current model. In each iteration, first the soft labels are fixed to update the model parameters; then the model parameters are fixed and the soft labels used in the next iteration are updated. This method achieved good performances for pathology image classification.
@@
He et al. [75] proposed a Self-Adaptation Network (SAN) using facial images to improve uncertain labeling for automatic depression estimation. Their architecture comprises: 1) an encoder for extracting deep features, 2) a self-attention module for learning the weights from the mini-batch and assigning lower weight to uncertain images, 3) a square ranking regularization module where the attention weights are ranked in descending order of importance and grouped into a high partition and a low partition, and 4) a re-label module for re-labeling the uncertain annotations for automatic depression estimation in the low partition. In the re-labeling module, for each image, the authors compute the softmax probabilities, which give the predicted probability distribution over all possible labels. For each image, the maximum value of the predicted probability is compared with the probability value of the provided label in the dataset. If the predicted probability is greater than the probability of the given label by a threshold value, the sample is assigned a new pseudo-label, which is based on the model’s prediction (the index of the maximum value of the predicted probability). Extensive experience on AVEC2013 and AVEC2014 demonstrated the efficiency of their proposed method.
@@
Zhu et al. [241] employed a hard sample aware self-training [^24] https://github.com/bupt-ai-cz/HSA-NRL/ strategy to correct and update labels. Their method comprises two phases: 1) a label correction and 2) a noise suppressing and hard enhancing scheme. The first phase produces "almost clean data" from noisy data. This is done by using an architecture consisting of 1) a classification/correction model, 2) an Easy/Hard/Noisy (EHN) detection scheme based on the sample training history and 3) a post-processing component. The main idea is to employ the prediction history for each sample as guiding information for differentiating hard from noisy samples. First, a classification network is trained on noisy data. The prediction history for one sample can then be stored for a specific number of epochs. The EHN detection module is then used to separate the data into easy, hard and noisy samples. In particular, the authors find that the normalized probability is easier to analyze than the loss value, a criteria used by previous state-of-the-art methods to detect clean samples. They propose to use the mean prediction value of the sample training history for detecting clean samples (clean samples have higher mean prediction probabilities than the noisy ones). They inject noise to the easy samples and retrain the classification model on the simulated noisy data. Based on the mean prediction probability value, they select hard and noisy samples. Next, a multi-layer perceptron classifier is trained on the training history of the initial classifier to distinguish noisy from hard samples. The correction model which is the same as the initial classification model is then trained on the easy and hard samples. After this step, the correction model is used to correct noisy labels. The labels of hard and noisy samples are corrected using the pseudo-labels produced by the correction model, which correspond to the class with the highest probability model output. These steps are repeated to further purify the dataset. Finally, in the post-processing step, the noisy data which can not be corrected are dropped out. In addition, hard samples whose labels were changed were also dropped out. The second phase employs a co-learning strategy based on co-teaching to train a noise robust model using the "almost clean data" (the second phase is detailed in Section 6.2.3). In Co-teaching, two DL networks are trained simultaneously, each DL network teach its peer network by selecting possibly clean data, data with small loss.
@@
Similar to Zhu et al. [241], Zhu et al. [242] proposed a co-teaching method to address noisy labels (see Section 6.2.3). Their framework includes a consistency-based noisy label correction module to detect noisy labels and correct them. It is a two-stage algorithm: 1) two networks are used to select clean samples according to their loss ranking, samples with the smallest loss are considered to be clean samples, 2) among the remaining suspected noisy data, samples that have consistent predictions on both networks are corrected. Specifically, they computed a prediction confidence score per network for the suspected noisy data. Samples are relabeled if they have the same predicted labels using both networks and the prediction confidences are larger than a predefined threshold. This means that if both networks agree on the classification of a sample and the confidence of this classification is high enough, the sample is assigned a new label based on this prediction. The new label (pseudo-label) is the class that both networks most strongly agree upon, under the condition that their prediction confidence surpasses the predetermined threshold. The rest uncorrected data are dropped and not used for training.
@@
Vindas et al. [198] proposed to estimate this transition matrix trough K-Nearest Neighbor (KNN) and local quality measures. Their approach involves four steps. First, features are extracted using an auto-encoder in an unsupervised manner. Second, they used t-distributed Stochastic Neighbor Embedding (t-SNE) algorithm, a dimensionality reduction technique, to project the features into a 2D space. In this step, the optimal projection is selected based on the silhouette score, which compares the similarity of a sample between the samples of its own class and the samples of the other classes. The best projection is based on the highest silhouette score. This is because for label propagation using KNN startegies, it is ideal to increase the inter-cluster distance and to reduce the intra-class distance; annotation errors mainly come from samples that are located at the boundaries of close class clusters, or from samples located in the wrong class cluster. Note that only labeled samples are used for this computation. Third, they perform automatic label propagation where the labels of high-quality labeled samples are propagated to high-quality unlabeled samples using KNN strategy and local quality metrics. This allows to increase the size of the training set. Fourth, for classification purposes, to compensate for the noise introduced by the automatic label propagation they used a robust loss function a generalized cross-entropy loss [233], formulated as follows:
@@
where $y_{i}$ and $f_{i}(x)$ are the i-th components of the true label $y$ and the predicted label $f(x)$, $v$ is a hyperparameter which allows control of the noise tolerance and the convergence speed; when $v\rightarrow 1$ we get the mean absolute error loss function whereas when $v\rightarrow 0$ we get the cross-entropy loss function. In their framework, $v$ was set to 0.7. They evaluated their framework on three tasks: emboli classification, organ classification and digit classification.
Similarly, Ying et al. [225] introduced label propagation for label recovery in the presence of noisy labels. They proposed a noisy label recovery algorithm based on Subset Label Iterative Propagation and Replacement (SLIPR) for dealing with noisy labels in COVID chest X-ray images classification. This algorithm aims to recover label and train the CNN on the label-recovered training set. It employs techniques such as principal component analysis, low-rank representation, and neighborhood graph regularization, followed by the kNN algorithm for classifying samples and recovering noisy labels through iterative label propagation and replacement. The first stage of their framework is a feature extraction and classification phase where they utilize a low-rank representation and a neighborhood graph regularization to extract both global and local features of the samples and KNN for classification purposes. The second stage consists of multi-level propagation and replacement of labels. In this stage, the concept of label propagation is used to select and replace the labels of the samples. In addition, they introduced a selection strategy for high confidence samples. Inspired by majority voting, the framework select high confidence samples as the training set based on the sample optional labels: a sample is considered to be high confidence sample if the majority result suggests that it should belong to the same type of label. For applying this principle, we should define the proportion of high confidence samples selected for training of the CNN, that is the minimum number of samples required for the classification model divided by the total number of samples. Based on this proportion, they followed these steps to select high confidence samples: 1) in each round of "label propagation and raplacement", they counted the optional labels of the samples 2) they ranked the confidence of the samples according to the majority voting principle, and 3) they selected the required number of high confidence samples.

### v3

@@ 5.1.3 Learning strategies
In order to reduce gradient variance from different domains, Atwany and Yaqub [66] presented Stochastic Weighted Domain Invariance [^18] https://github.com/BioMedIA-MBZUAI/DRGen, a method leveraging the Fishr regularization coupled with iteration-wise avergaging of weights (SWA). It is built upon Stochastic Weight Averaging Densely (SWAD) [139] and Fishr [140] to encourage seeking a flatter minima while imposing a regularization. SWAD seeks a flat minimum by averaging the weights by iterations (rather than by epochs). It enables averaging weights only from specific iterations where the validation loss decreases. On the other hand, Fishr [140] is a regularization approach that enforces domain invariance in the space of the gradients of the loss. In particular, the domain-level variances of gradients are matched across training domains. Fishr regularization enforces the domain-level gradient invariance in the classifier by aligning the gradient covariances at the domain level. The Fishr loss is thus formulated as follows:
@@
Bissoto et al. [59] proposed a DG approach based on DRO. Their pipeline involved first partitioning the data into training and test sets, with amplified correlations between artifacts and class labels (malignant vs. benign), which appear in opposite directions in the dataset splits. The training set was then divided into artifact-based domains. Next, the GroupDRO [141] algorithm, which minimizes the loss of the worst-case training source environment, was then trained on these artifact-based domains. In the last phase, the authors employed a test-time debiasing procedure to reduce the influence of spurious features in the inference images. The experimental results for skin lesion detection showed that GroupDRO allows learning more robust features.
@@
Lin et al. [67] proposed a multi-task network [^19] https://github.com/linzhlalala/CVD-risk-based-on-retinal-fundus-images for cardiovascular disease risk (CVD) estimation using fundus photographs. To learn invariant representations, a Siamese network was pre-trained using the left and right fundus photographs for each patient as positive sample pairs. This network was then jointly trained on WHO-CVD score and on seven clinical variables explicitly correlated with WHO-CVD such as age, systolic blood pressure and gender. They also integrated a feature-level knowledge distillation. Given two input images (left and right fundus photographs) from a single patient, the feature with the smallest supervised learning loss is considered as the teacher whereas the other as the student. For the teacher-level features, they performed stop-gradient operation when updating the feature extractor. The results showed that the pre-training strategy reduced the feature-space discrepancy between the UK biobank dataset collected using the Topcon 3D OCT-1000 MKII and the other cameras (Mediwork portable camera).
@@
Wang et al. [68] used the same approach for UDA by integrating auxiliary task (predicting age, gender and race) to their framework. Their method consisted of four stages: 1) pre-training a classifier with a feature extractor, an auxiliary task network and a primary task network using source data, 2) fine-tuning the feature extractor and the auxiliary task network for the auxiliary tasks using target data while constraining the feature extractor not to change significantly, 3) fine-tuning the primary classifier on the source data to correctly classify the primary task (i.e., classification) based on the modified features, 4) performing inference using updated weights on the target data. The authors showed improvement of performances when tested on 3D brain MRI dataset for classifying Alzheimer’s disease and schizophrenia.
@@
For MIDOG challenge, Razavi et al. [69] proposed a multi-stage mitosis detection method based on a Cascade R-CNN. The Cascade-RCNN comprises a sequential detectors with increasing intersection over union to reduce false positives. It consists of two-stage: 1) a region proposal network that detects candidate region and 2) a region proposal network and a classification network that performs classification on the candidate regions. This method achieved a F1 score of 0.7492 on the MIDOG testing set. This ends our presentation of DG solutions to address covariate shift.
5.2 Concept shift and label noise in medical image classification
@@
The quality of annotations has a crucial role on the model generalizability. Nevertheless, the annotation process can be subjective, and the issue of label noise is sometimes unavoidable. In addition, the quality of annotations can differ among various annotators. To improve prediction performances, DL methods have been proposed to address the problem of noisy labels and concept shift. These methods can be categorized into two main classes: data-centric methods and model-centric methods. Data-centric methods focus on data adjustment and transformation (Section 5.2.1). These methods focus on identifying noisy samples and correcting their labels. Model-centric methods include learning methods (Section 5.2.2) and collaborative methods (Section 5.2.3). Learning methods propose an optimization framework based on loss functions (i.e., regularization), architecture (e.g., graphs) and learning strategies (active learning, zero-shot learning, meta-learning, etc.). On the other hand, collaborative methods exploit the cooperation between models to boost DL performances.
@@ 5.2.1 Data adjustment and transformation
For instance, Son et al. [70] proposed to detect mislabeled samples based on the classifier’s confidence (i.e., the softmax outputs) and to mask the loss function computed over such samples [^20] https://bitbucket.org/woalsdnd/codes_and_data. Noisy labels were simulated in the training data by randomly flipping labels with probabilities ranging between 0 and 0.8. To detect mislabeled samples, a filtration network was trained on top of the classification network to minimize the logistic regression loss over validation data (i.e., clean data). It examines the training set with positive labels and assigns high values on positive images with clean labels and low values to suspected negative images. For the task of referable DR detection, this method outperformed state-of-the art methods such as S-model and Bootstrap at noise ratios of 0.2 and 0.4 on the Kaggle 2015 dataset. In addition, integrating the classifier’s confidence in an active learning scheme showed good results ranking first on the PALM challenge with an AUC of 0.9993 for pathological myopia classification.
@@
Xue et al. [71] used confident learning [142] to identify noisy samples in the training data. It employs the predicted probability outputs (self-confidence) and the noisy labels for estimating label uncertainty, i.e., the joint distribution between the noisy and true labels. The class imbalance and heterogeneity in predicted probability distributions across classes are addressed by using a per-class threshold (expected self-confidence for each class) when calculating the confident joint. To further enhance the precision, Xue et al. [71] proposed an ensemble strategy consisting of training three different classification networks and selecting the candidates that were jointly identified as noisy using confident learning. For automated visual evaluation for precancer screening, they achieved a kappa score of 0.687 with the cleaned development set, compared to 0.682 with the original noisy development set.
Aljuhani et al. [72] presented Uncertainty-Aware Sampling Framework (UASF) [^21] https://github.com/machiraju-lab/UA-CNN to tackle the problem of weak labels in digital pathology, where WSI-level diagnoses lack precise annotations indicating specific regions within WSI responsible for the diagnosed label. This method employs an informative sampling algorithm to select the most relevant tiles by estimating uncertainties using variational Monte Carlo inference, with the predictive entropy as a measure. The relevant tiles are identified by their high prediction probability and low uncertainty. Once the disease-representative tiles were effectively identified, the prediction performance was enhanced by training the model on the refined training dataset. For the leiomyosarcoma histological subtype grading task, this approach achieved 83% accuracy.
@@
Bai et al. [73] adopted a convolutional bootstrapping strategy [^22] https://github.com/BRF123/Cnngeno to handle noisy labels in the data. First, a set of highly reliable seeds (i.e., a subset of samples) were manually selected as training set and the model was trained until convergence. The model was then used to classify the remaining samples in the dataset. The process involves selecting samples with higher classification confidence into a seed set, based on a classification confidence level set at 0.8. The expansion of the seed set is repeated until no new seeds are added, and the final trained model is then used for classification. For the calling of structural variation genotype, the proposed method performed better than the current state-of-the-art methods on complex real data with high and low coverage.
Xu and Chen [74], Hu et al. [75] proposed a sample re-weighting algorithm that assigns weights to training samples, with higher weights assigned to clean samples and lower weights to noisy samples. In [74], these weights are determined to minimize the loss on a clean unbiased validation dataset. The authors showed good performances when evaluating their method on calcium imaging data of anterior lateral motor cortex, with an F1 score and balanced accuracy greater than 0.85, despite noise levels varying between 9% to 52%. In contrast to this method, Hu et al. [75] [^23] https://github.com/TwistedW/MIAV used the concept of sample interaction in small groups as in Peng et al. [143] which does not require a clean validation set. For the automated classification of retinal arteries and veins, they achieved an accuracy of 97.47%, 96.91%, 97.79%, and 98.18% on AV-DRIVE, HRF, LES-AV and a private dataset, respectively.
@@
To leverage incomplete observations, Hermoza et al. [76] [^24] https://github.com/renato145/CASurv used the concept of pseudo labels, where the output of the network is used to estimate the label. The authors argue that the quality of generated pseudo-labels depends on the training procedure stage: during the first epochs, the pseudo-labels are less accurate than those at the last epochs. To address this issue, they used a cosine annealing schedule to control the generation of pseudo-labels during training. The evaluation of their proposed method on pathology and X-ray images from the TCGAGM and NLST datasets showed good prediction survival accuracy on both datasets.
@@
Inspired by epistemic uncertainty [144], Bai et al. [77] proposed Pseudo-Labeling based on Adaptive Threshold (PLAT) to reduce the generation of noisy labels in a semi-supervised approach. Unlabeled images are inputted to the model $k$ times using Monte Carlo Dropout, resulting in $k$ predictions. Uncertainty of pixels is estimated by computing the variance of these predictions, and then normalized by dividing by the largest variance among all predictions. The normalized result is used as an adaptive threshold. Compared to model trained only on labeled images, this method showed a gain of 9%-13% in terms of F1 score.
@@
Qiu et al. [78] proposed a self-training strategy consisting of noisy label cleaning optimization. Initially, the model is pretrained with the noisy labels using a large fixed learning rate, under the assumption that the network can avoid overfitting to the label noise. Then, noise-free labels (soft labels) are computed using the softmax output of the pretrained model. During each iteration, the soft labels are fixed to update the model parameters; then the model parameters are fixed, and the soft labels are updated for the next iteration. This method achieved good performances for pathology image classification.
@@
He et al. [79] proposed a re-labeling module in their Self-Adaptation Network (SAN). For each image, they computed the softmax probabilities. Then, the maximum value of the predicted probability is compared with the probability value of the provided label in the dataset. If the predicted probability is greater than the probability of the given label by a threshold value, the sample is assigned a new pseudo-label, which is based on the model’s prediction. Extensive experience on AVEC2013 and AVEC2014 demonstrated the efficiency of their proposed method for automatic depression detection.
@@
Zhu et al. [80] employed a hard sample aware self-training [^25] https://github.com/bupt-ai-cz/HSA-NRL/ strategy to correct and update labels. They used the mean prediction value of the sample training history of their classification model to separate the data into easy, hard and noisy samples. Their classification network was first trained on noisy data. Easy samples were identified as those with have higher mean prediction probabilities. After selecting the easy samples, noisy samples were simulated by injecting noise to the easy samples and the classification model was retrained on the simulated noisy data. Based on the mean prediction probability value, clean samples of the noisy data were identified and the rest was utilized for training a multi-layer perceptron classifier, which was designed to distinguish between hard and noisy samples using the training history of the initial classifier as input. The classification model was then retrained on the easy and hard samples. After this step, the labels of hard and noisy samples are corrected using the pseudo-labels produced by the classification model, which correspond to the class with the highest probability model output. These steps are repeated to further purify the dataset. Finally, in the post-processing step, the noisy data with unchanged labels and the hard samples with changed labels were dropped out.
@@
Zhu et al. [81] included a consistency-based noisy label correction module in their framework to detect noisy labels and correct them. It is a two-stage algorithm: 1) two networks are used to select clean samples according to their loss ranking, samples with the smallest loss are considered to be clean samples, 2) among the remaining suspected noisy data, samples that have consistent predictions on both networks are corrected. The new label (pseudo-label) is assigned as the class that both networks most strongly agree upon, under the condition that their prediction confidence surpasses the predetermined threshold.
@@
Vindas et al. [83] proposed to estimate this transition matrix trough K-Nearest Neighbor (KNN) and local quality measures. Their approach involves four steps. First, features are extracted using an auto-encoder in an unsupervised manner. Second, t-distributed Stochastic Neighbor Embedding (t-SNE) algorithm was used to project the features into a 2D space. In this step, the optimal projection was selected based on the silhouette score. Note that only labeled samples are used for this computation. Third, the labels of high-quality labeled samples were propagated to high-quality unlabeled samples using KNN strategy and local quality metrics. This allows to increase the size of the training set. Fourth, for classification purposes, to compensate for the noise introduced by the automatic label propagation, a robust loss function a generalized cross-entropy loss [146], was introduced as follows:
@@
where $y_{i}$ and $f_{i}(x)$ are the i-th components of the true label $y$ and the predicted label $f(x)$, $v$ is a hyperparameter which allows control of the noise tolerance and the convergence speed; when $v\rightarrow 1$ we get the mean absolute error loss function whereas when $v\rightarrow 0$ we get the cross-entropy loss function. In their framework, $v$ was set to 0.7. This framework was evaluated on three tasks: emboli classification, organ classification and digit classification.
Ying et al. [84] proposed a noisy label recovery algorithm based on Subset Label Iterative Propagation and Replacement (SLIPR) for dealing with noisy labels in COVID chest X-ray images classification. This algorithm aims to recover label and train the CNN on the label-recovered training set. The first stage of their framework is a feature extraction and classification phase where they utilize a low-rank representation and a neighborhood graph regularization to extract both global and local features of the samples and KNN for classification purposes. The second stage consists of multi-level propagation and replacement of labels. In this stage, the concept of label propagation is used to select and replace the labels of the samples. In addition, a selection strategy for high confidence samples was introduced. Inspired by majority voting, it selects high confidence samples as the training set based on the sample optional labels: a sample is considered to be high confidence sample if the majority result suggests that it should belong to the same type of label.